anthony122:

For a right circular cone, if the radius is reduced to 1/5 its original size and the slant height is reduced to 1/6 its original size, find the surface area if the original radius is 8 centimeters and the original slant height is 13 centimeters. *** edited by a moderator to supply missing information ***

jhonyy9:

do you can reposting these jpg files using the Attach File button bottom to this window hope will be more understandably

Vocaloid:

I have added the original numbers back in if you would like to help

Angle:

Here is a cool source you can use and play with the graphics they provide. https://www.mathopenref.com/conearea.html The Area of a Right Cone = \(\large ( \pi *r*s ) + ( \pi *r^2) \) where r = radius and s = slant height Now, we are told that the original radius is 8 cm and original slant height is 13 cm. We are also told that: the new radius is 1/5 the original which makes the new radius = 8/5 cm and the new slant height is 1/6 the original which makes the new slant height = 13/6 cm Using these two new values, we can plug them into the provided equation (above) and find the answer to this problem. :)